Rosa's age is twice Rod's age and Lily's age is 5 years younger than Rosa. The sum of Rosa's and Rod's age is 3 less than twice Lily's age. Find the ages of the three.
Answer by mananth(16946)

let Rod's age = x
Rosa's age = 2x
Lily's age = 2x-5
...
2x+x=2(2x-5)-3
3x=4x-10-3
3x-4x=-13
-x=-13
x= 13 Rod's age
Rosa = 13*2 = 26 years
Lily = 2*13-5 = 21 years
...
m.ananth@hotmail.ca